# Predicting Subway Passenger Flows under Incident Situation with Causality

### v1

@@ 4.5 Validation of sample selection
In the method section, we proposed that it is important to select samples that are significantly impacted to calibrate causal effect prediction model. In deployment, choosing passenger flows which are significantly affected by incidents is also beneficial. To validate this statement, and to analyze how different threshold values ($p1$ and $p2$ in Figure 4) affect overall prediction error, we conducted further analysis.
@@ 5 Conclusion
There are several limitations in our current work which could be improved. First, our dataset is relatively small, covering only three months and six serious incidents. This limited sample size may affect the model’s robustness, suggesting the need for a larger dataset in future studies.
Second, the current causal identification and causal effect prediction models rely on traditional methods, without accounting for potential spatial relationships in causal effects. In the future, spatial information could be used to estimate causal effects more accurately. For instance, in estimating counterfactual passenger flow through synthetic control methods, it might be helpful to incorporate spatially related flows to decide weights W in Eq3. Additionally, traditional models, such as Random Forest and linear regression, are used as causal effect prediction models in this research. In the future, using deep spatial-temporal neural networks might lead to improvements. Moreover, incorporating more features, such as bus schedules and land usage, may improve the prediction results.
Finally, forecasting passenger flow under special conditions remains challenging due to limited data on such scenarios. When collecting sufficient data proves difficult, large language models (LLMs) could offer an alternative, as recent research has shown that LLMs, with their extensive world knowledge and generalization ability, can make reasonable predictions for unusual scenarios [13, 17]. Integrating LLMs for subway passenger flow prediction under incident conditions may be a promising future direction.
@@
$E(\beta-\hat{\beta})^{2}=(1-p)\ |\beta|^{2}+c\frac{(p\sigma_{2}^{2}+\sigma_{1}^{2})}{n}$
@@
$E\left(\beta-\hat{\beta}\right)^{2}$ $=E\left(\beta-E\hat{\beta}+E\hat{\beta}-\hat{\beta}\right)^{2}$
$=E\left(\beta-E\hat{\beta}\right)^{2}+E\left(E\hat{\beta}-\hat{\beta}\right)^{2}+2E\left[\left(\beta-E\hat{\beta}\right)\left(E\hat{\beta}-\hat{\beta}\right)\right]$
@@
$E\left(\beta-\hat{\beta}\right)^{2}=E\left(\beta-E\hat{\beta}\right)^{2}+E\left(E\hat{\beta}-\hat{\beta}\right)^{2}=bias\left(\hat{\beta}\right)^{2}+Tr(Var\left(\hat{\beta}\right))$
@@
$E\left(\hat{\beta}\right)=E\left[\left(XX^{T}\right)^{-1}X^{T}Y\right]=\left(XX^{T}\right)^{-1}X^{T}E\left(Y\middle|X\right)$
@@
$E\left(Y\middle|X\right)=(1-p)\times E\epsilon_{1}+pE\left(\epsilon_{1}+\left(\beta^{T}X+\epsilon_{2}\right)\right)=pX\beta$
@@
$E\left(\hat{\beta}\right)=E\left(1-p\right)\left(XX^{T}\right)^{-1}X^{T}X\beta=p\beta$
@@
$bias\left(\hat{\beta}\right)^{2}=\left((1-p)\beta\right)^{2}$
@@
$Var\left(\hat{\beta}\right)=Var\left(\left(XX^{T}\right)^{-1}X^{T}Y\right)=\left(XX^{T}\right)^{-1}X^{T}Var\left(Y\middle|X\right)X\left(XX^{T}\right)^{-1}$
@@
$Var\left(Y\middle|X\right)=\left((1-p)\sigma_{1}^{2}+p\left(\sigma_{1}^{2}+\sigma_{2}^{2}\right)\right)I_{n}=\left(\sigma_{1}^{2}+p\sigma_{2}^{2}\right)I_{n}$
Where $I_{n}$ is $n\times n$ identical matrix. Substituting it into $Var\left(\hat{\beta}\right)$
$Var\left(\hat{\beta}\right)=\left(XX^{T}\right)^{-1}X^{T}\left(\sigma_{1}^{2}+p\sigma_{2}^{2}\right)I_{n}X\left(XX^{T}\right)^{-1}$
@@
$Tr\left(Var(\hat{\beta})\right)=\left(\sigma_{1}^{2}+p\sigma_{2}^{2}\right)Tr(\left(XX^{T}\right)^{-1})$
Suppose each entry in $X$ is i.i.d sample from $N\left(0,\sigma_{x}^{2}\right)$, then $XX^{T}$ is Wishart distribution [4] and:
$E{Tr\left(XX^{T}\right)}^{-1}=\frac{d}{(n-d-1)\sigma_{x}^{2}}$
@@
$Tr\left(Var\left(\hat{\beta}\right)\right)=\left(\sigma_{1}^{2}+p\sigma_{2}^{2}\right)\frac{d}{(n-d-1)\sigma_{x}^{2}}$
@@
$Tr\left(Var\left(\hat{\beta}\right)\right)\approx\left(\sigma_{1}^{2}+p\sigma_{2}^{2}\right)\frac{d}{n\sigma_{x}^{2}}$
@@
$E\left(\beta-\hat{\beta}\right)^{2}={(1-p)}^{2}\beta^{2}+\left(\sigma_{1}^{2}+p\sigma_{2}^{2}\right)\frac{d}{n\sigma_{x}^{2}}={(1-p)}^{2}\beta^{2}+c\frac{\sigma_{1}^{2}+p\sigma_{2}^{2}}{n}$
where:
$c=\frac{d}{\sigma_{x}^{2}}$
$\qed$
@@
$E\left(y-\hat{y}\right)^{2}=\frac{1}{2}P^{2}(Ef\left(x\right)^{2}+E\hat{f}\left(x\right)^{2}-E\left(\hat{f}\left(x\right)-f\left(x\right)\right)^{2})-PE\hat{f}\left(x\right)^{2}+c$
@@
[…]
@@
$E\left(y-\hat{y}\right)^{2}=$ $\;E\Big{[}\int_{P}^{1}t\left(\hat{f}(x)-\left(\epsilon_{1}+\left(f(x)+\epsilon_{2}\right)\right)\right)^{2}dt+\int_{P}^{1}(1-t)\left(\hat{f}(x)-\epsilon_{1}\right)^{2}dt$
$+\int_{0}^{P}t\left(\epsilon_{1}+\left(f(x)+\epsilon_{2}\right)\right)^{2}dt+\int_{0}^{P}(1-t)\epsilon_{1}^{2}dt\Big{]}.$
@@
$E\left(y-\hat{y}\right)^{2}=$ $\;\frac{1}{2}P^{2}\left(Ef\left(x\right)^{2}+E\hat{f}\left(x\right)^{2}-E\left(\hat{f}\left(x\right)-f\left(x\right)\right)^{2}\right)$
$-PE\hat{f}\left(x\right)^{2}+E\left(\hat{f}(x)^{2}+\frac{f\left(x\right)^{2}}{2}-f\left(x\right)\hat{f}\left(x\right)\right)+\sigma_{1}^{2}+\frac{\sigma_{2}^{2}}{2}$
$=$ $\;\frac{1}{2}P^{2}\left(Ef\left(x\right)^{2}+E\hat{f}\left(x\right)^{2}-E\left(\hat{f}\left(x\right)-f\left(x\right)\right)^{2}\right)-PE\hat{f}\left(x\right)^{2}+c.$

### v2

@@ 4.5 Validation of sample selection
In the method section, we proposed that it is important to select samples that are significantly impacted to calibrate causal effect prediction model. In deployment, choosing passenger flows that are significantly affected by incidents is also beneficial. To validate this statement, and to analyze how different threshold values ($p1$ and $p2$ in Figure 4) affect overall prediction error, we conducted further analysis.
@@ 5 Conclusion
There are several limitations in our current work which could be improved. First, the current 30-minute prediction interval introduces latency in emergency response. Future work should develop sparse-data tolerant models (e.g., zero-inflated regression) and take use of larger dataset for finer-grained prediction. Besides, the reliance on post-incident features (e.g., maximum delay time) impedes real-time deployment. Subsequent studies need to use immediately available incident descriptors (e.g., incident severity level, incident type) for causal effect estimation. Moreover, our model treats ODs as independent units, neglecting network-wide behavioral interdependencies. Collecting larger-scale incident datasets would enable spatiotemporal graph-based approaches to capture passenger rerouting dynamics.
Another major limitation lies in the dataset: it only contains six eligible incidents—primarily caused by signal failures, mechanical and electrical equipment malfunctions. However, during extreme events like fires or terrorist attacks, passenger behavioral responses would diverge significantly from those observed in our dataset. This restricts our model’s generalizability. Future studies with datasets encompassing diverse incident types may address this by calibrating separate models for distinct emergency categories.
Finally, forecasting passenger flow under special conditions remains challenging due to limited data on such scenarios. When collecting sufficient data proves difficult, large language models (LLMs) could offer an alternative, as recent research has shown that LLMs, with their extensive world knowledge and generalization ability, can make reasonable predictions for unusual scenarios [Huang2024EnhancingTP, liang2024exploring]. Integrating LLMs for subway passenger flow prediction under incident conditions may be a promising future direction.
@@
Author contributions
Conceptualisation: Chao Yang, data curation: Xiannan Huang, methodology: Chao Yang+Xiannan Huang+Shuhan Qiu, software: Xiannan Huang, formal analysis: Xiannan Huang, investigation: Xiannan Huang+Quan Yuan, writing-original draft: Xiannan Huang, writing-review and editing: Shuhan Qiu+Quan Yuan+Chao Yang, visualisation: Xiannan Huang, supervision: Chao Yang. All authors reviewed the results and approved the final version of the manuscript.
Statements and Declarations
Conflict of interest
The authors declare no competing interests.
@@
$E(\beta-\hat{\beta})^{2}=(1-p)\ |\beta|^{2}+c\frac{(p\sigma_{2}^{2}+\sigma_{1}^{2})}{n}$ (15)
@@
$E\left(\beta-\hat{\beta}\right)^{2}$ $=E\left(\beta-E\hat{\beta}+E\hat{\beta}-\hat{\beta}\right)^{2}$ (16)
[…]
@@
$)E\left(\beta-\hat{\beta}\right)^{2}=E\left(\beta-E\hat{\beta}\right)^{2}+E\left(E\hat{\beta}-\hat{\beta}\right)^{2}=bias\left(\hat{\beta}\right)^{2}+Tr(Var\left(\hat{\beta}\right$)) (18)
@@
$E\left(\hat{\beta}\right)=E\left[\left(XX^{T}\right)^{-1}X^{T}Y\right]=\left(XX^{T}\right)^{-1}X^{T}E\left(Y\middle|X\right)$ (19)
@@
$E\left(Y\middle|X\right)=(1-p)\times E\epsilon_{1}+pE\left(\epsilon_{1}+\left(\beta^{T}X+\epsilon_{2}\right)\right)=pX\beta$ (20)
@@
$E\left(\hat{\beta}\right)=E\left(1-p\right)\left(XX^{T}\right)^{-1}X^{T}X\beta=p\beta$ (21)
@@
$bias\left(\hat{\beta}\right)^{2}=\left((1-p)\beta\right)^{2}$ (22)
@@
$Var\left(\hat{\beta}\right)=Var\left(\left(XX^{T}\right)^{-1}X^{T}Y\right)=\left(XX^{T}\right)^{-1}X^{T}Var\left(Y\middle|X\right)X\left(XX^{T}\right)^{-1}$ (23)
@@
$Var\left(Y\middle|X\right)=\left((1-p)\sigma_{1}^{2}+p\left(\sigma_{1}^{2}+\sigma_{2}^{2}\right)\right)I_{n}=\left(\sigma_{1}^{2}+p\sigma_{2}^{2}\right)I_{n}$ (24)
[…]
$Var\left(\hat{\beta}\right)=\left(XX^{T}\right)^{-1}X^{T}\left(\sigma_{1}^{2}+p\sigma_{2}^{2}\right)I_{n}X\left(XX^{T}\right)^{-1}$ (25)
@@
$Tr\left(Var(\hat{\beta})\right)=\left(\sigma_{1}^{2}+p\sigma_{2}^{2}\right)Tr(\left(XX^{T}\right)^{-1})$ (26)
Suppose each entry in $X$ is i.i.d sample from $N\left(0,\sigma_{x}^{2}\right)$, then $XX^{T}$ is Wishart distribution [bilodeau1999theory] and:
$E{Tr\left(XX^{T}\right)}^{-1}=\frac{d}{(n-d-1)\sigma_{x}^{2}}$ (27)
@@
$Tr\left(Var\left(\hat{\beta}\right)\right)=\left(\sigma_{1}^{2}+p\sigma_{2}^{2}\right)\frac{d}{(n-d-1)\sigma_{x}^{2}}$ (28)
@@
$Tr\left(Var\left(\hat{\beta}\right)\right)\approx\left(\sigma_{1}^{2}+p\sigma_{2}^{2}\right)\frac{d}{n\sigma_{x}^{2}}$ (29)
@@
$E\left(\beta-\hat{\beta}\right)^{2}={(1-p)}^{2}\beta^{2}+\left(\sigma_{1}^{2}+p\sigma_{2}^{2}\right)\frac{d}{n\sigma_{x}^{2}}={(1-p)}^{2}\beta^{2}+c\frac{\sigma_{1}^{2}+p\sigma_{2}^{2}}{n}$ (30)
where: $c=\frac{d}{\sigma_{x}^{2}}$
@@
$E\left(y-\hat{y}\right)^{2}=\frac{1}{2}P^{2}(Ef\left(x\right)^{2}+E\hat{f}\left(x\right)^{2}-E\left(\hat{f}\left(x\right)-f\left(x\right)\right)^{2})-PE\hat{f}\left(x\right)^{2}+c$ (31)
@@
$E\left(y-\hat{y}\right)^{2}=$ $\;E\int P\left(p_{i}>P\ \text{and}\ y_{i}=\epsilon_{1}+\left(f(x_{i})+\epsilon_{2}\right)\right)$ (32)
$\times\left(\hat{f}(x_{i})-\left(\epsilon_{1}+\left(f(x_{i})+\epsilon_{2}\right)\right)\right)^{2}\,dp_{i}$ (33)
$+\int P\left(p_{i}>P\ \text{and}\ y_{i}=\epsilon_{1}\right)\left(\hat{f}(x_{i})-\epsilon_{1}\right)^{2}\,dp_{i}$ (34)
$+\int P\left(p_{i}<P\ \text{and}\ y_{i}=\epsilon_{1}\right)\epsilon_{1}^{2}\,dp_{i}$ (35)
$+\int P\left(p_{i}<P\ \text{and}\ y_{i}=\epsilon_{1}+\left(f(x_{i})+\epsilon_{2}\right)\right)\left(\epsilon_{1}+\left(f(x_{i})+\epsilon_{2}\right)\right)^{2}\,dp_{i}$ (36)
@@
$E\left(y-\hat{y}\right)^{2}=$ $\;E\Big[\int_{P}^{1}t\left(\hat{f}(x)-\left(\epsilon_{1}+\left(f(x)+\epsilon_{2}\right)\right)\right)^{2}dt+\int_{P}^{1}(1-t)\left(\hat{f}(x)-\epsilon_{1}\right)^{2}dt$ (37)
$+\int_{0}^{P}t\left(\epsilon_{1}+\left(f(x)+\epsilon_{2}\right)\right)^{2}dt+\int_{0}^{P}(1-t)\epsilon_{1}^{2}dt\Big]$ (38)
@@
$E\left(y-\hat{y}\right)^{2}=$ $\;\frac{1}{2}P^{2}\left(Ef\left(x\right)^{2}+E\hat{f}\left(x\right)^{2}-E\left(\hat{f}\left(x\right)-f\left(x\right)\right)^{2}\right)$ (39)
$-PE\hat{f}\left(x\right)^{2}+E\left(\hat{f}(x)^{2}+\frac{f\left(x\right)^{2}}{2}-f\left(x\right)\hat{f}\left(x\right)\right)+\sigma_{1}^{2}+\frac{\sigma_{2}^{2}}{2}$ (40)
$=$ $\;\frac{1}{2}P^{2}\left(Ef\left(x\right)^{2}+E\hat{f}\left(x\right)^{2}-E\left(\hat{f}\left(x\right)-f\left(x\right)\right)^{2}\right)-PE\hat{f}\left(x\right)^{2}+c$ (41)